OpenStudy (sasogeek):

when you divide 2 terms with variables, each with coefficients in parenthesis, and outside the parenthesis, both terms have same exponents but the coefficients are different, do you divide the coefficients and make the variables the same and subtract the exponents?

OpenStudy (sasogeek):

eg. (20x)^2 / (5x)^2 =400x^2 / 25x^2=16x^0 ?

hartnn (hartnn):

thats correct and equal to 16.

OpenStudy (sasogeek):

ok thanks :)

hartnn (hartnn):

or u can do like this: \(\large\frac{(20x)^2}{(5x)^2}=(20x/5x)^2=4^2=16\)

OpenStudy (sasogeek):

ok thanks, interesting concept xD

hartnn (hartnn):

welcome :)